OpenStudy (anonymous):

If the rate of disappearance H2 = -0.13 M/s at a specific time, what is the value for the rate of formation of HCl during the same interval?

OpenStudy (xishem):

This relationships depends on the mechanism. What is the chemical equation of the reaction you are analyzing?

OpenStudy (anonymous):

H2 + 2 ICl  I2 + 2HCl…

OpenStudy (anonymous):

Plz help me understand the concept

OpenStudy (anonymous):

how will I find the value of rate?

OpenStudy (xishem):

Let's forget about rate for a second. If one mol of H2 has been reacted, how many moles of HCl is formed?

OpenStudy (anonymous):

I dont know:(

OpenStudy (anonymous):

2

OpenStudy (anonymous):

as we can see in the product side

OpenStudy (anonymous):

2HCL

OpenStudy (xishem):

That's right. Based on the coefficients, you know that each molecule (and mol) of H2 will go to form two molecules (and mol) HCl.

OpenStudy (anonymous):

yes

OpenStudy (xishem):

So we can make an assumption that over any time interval, the ratio of H2 consumed to the ratio of HCl produced will be... \[H_{2_{consumed}}:HCl_{produced}=1:2\]

OpenStudy (xishem):

Let's look back at the original problem now.

OpenStudy (anonymous):

yes 1:2

OpenStudy (xishem):

If the rate of disappearance of H2 is -0.13M/s:\[Rate_{H_2}=-0.13\frac{M}{s} \]That means that every second, one M of H2 is being reacted. What does 1M mean in terms of moles?

OpenStudy (anonymous):

the molecule of Hydrogen

OpenStudy (anonymous):

that is disappeared every second and making two molecules of HCL

OpenStudy (xishem):

So if you have twice the number of molecules disappearing over the same time interval, what is the rate of disappearance of HCl?

OpenStudy (anonymous):

1 mol

OpenStudy (xishem):

\[Rate=\frac{molecules}{t}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

whre as - shows the decrease in the rate of molecules

OpenStudy (xishem):

Where delta A is the number of molecules?

OpenStudy (xishem):

*Change in number of molecules.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this the right equation?

OpenStudy (anonymous):

what is the time in this case?

OpenStudy (xishem):

The rate must be multiplied by the reciprocal of the stoichiometric coefficient to account for what we just discussed. For a chemical equation:\[aA+bB \rightarrow cC+dD\]Where lower case letters are stoichiometric coefficients and capital letters are some chemical species, the following is true:\[Rate_{rxn}=\frac{-1}{a}\frac{\Delta [A]}{\Delta t}=\frac{-1}{b}\frac{\Delta [B]}{\Delta t}=\frac{1}{c}\frac{\Delta [C]}{\Delta t}=\frac{1}{d}\frac{\Delta [D]}{\Delta t}\]Does that make sense? This is basically the equation you gave, except it didn't take into account the stoichiometric coefficients.

OpenStudy (anonymous):

I got it . but in this question how will I find the value ?

OpenStudy (anonymous):

do I have to add the values of all the reactant and product sides/